Churchill Bank has reserves of $100mn and loans to clients of $500mn. In addition, it has securities assets which act as collateral for the bank's own borrowing of $280mn (assume a 30% haircut was taken on the assets). The bank's only other debt is customer deposits of $500mn.

a. draw up a balance sheet for Churchill Bank

b. what is the bank's leverage ratio?

c. what is the bank's gearing?

d. what is the bank's reserves ratio?

e. what happens to the leverage multiple if the accounting value of the bank's loans is estimated to have fallen by 40%?
Answer #1

A) Balance sheet

Asset -

Receivables - $ 500 mm

Securities - $ 196 mm (after haircut adjustment)

Liabilities + Capital

Reserves - $ 100 mm

Payables - $ 500 mm

Equity - $ 96 mm

B) Leverage ratio = Total debt/Shareholder's equty = 500/(96+100) = 2.55

C) Borrowed funds/Equity = 500/196 = 2.55